Find the molar concentration of the sulfuric acid from the titration result:

Complete the calculation given below:

No. of moles of KOH in 25.00 cm3

=





=

………………………… (1 mark)

\ no. of moles H2SO4 in 15.00 cm3

=





=

………………………… (1 mark)

\ molar concentration of H2SO4

=



(Include units in your answer)

=

………………………… (2 marks)
1
Expert's answer
2021-07-18T01:28:25-0400

Relative Molecular Mass (RMM) of

KOH is 56.11 g/mole


No. of moles of KOH in 25.00 cm3"= \frac{0.95}{56.11}=0.0169moles"


2KOH (aq) + H2SO4 (aq) → K2SO4 (aq) + 2H2O (l)


Moles of "H_2SO_4" "=\frac{0.0169}{2}=0.00845moles"


Molarity ="\frac{0.00845}{0.015}=0.563M"